# Spatial distribution of organics in the Horsehead nebula: signposts of chemistry driven by atomic carbon

### v2

@@
Complex organic molecules (COMs) are considered essential precursors to prebiotic species in the interstellar and circumstellar medium. Despite their astrobiological relevance, many aspects of COMs formation remain unclear, particularly the role of ultraviolet (UV) radiation. While COMs were once expected to be efficiently destroyed under UV-irradiated conditions, detections in photodissociation regions (PDRs) have challenged this view. However, the mechanisms by which UV radiation contributes to their formation are still uncertain. Here, we present moderately resolved maps of simple and complex organic molecules at the UV-illuminated edge of the Horsehead nebula, obtained by combining Atacama Large Millimeter/submillimeter Array (ALMA) and IRAM 30m single-dish observations at $\sim$$15^{\prime\prime}$ resolution. For the first time in this PDR environment, we analyze the spatial distribution of species such as C17O, CH2CO, CH3CHO, HNCO, CH3CN, and HC3N. By incorporating previous C17O and C18O single-dish data as well as Plateau de Bure Interferometer (PdBI) maps of H2CO and CH3OH, we derive profiles of gas density, temperature, thermal pressure, and column densities of the organic species as a function of distance from the UV source. Our results show that most organic species—particularly H2CO, CH2CO, CH3CHO, HNCO, and CH3CN—exhibit enhanced column densities at the UV-illuminated edge compared to cloud interiors, possibly indicating efficient dust-grain surface chemistry driven by the diffusion of atomic C and radicals produced via photodissociation of CO and CH3OH, as supported by recent laboratory experiments. The exceptions, HC3N and CH3OH, can be attributed to inefficient formation on dust grains and ineffective non-thermal desorption into the gas phase, respectively. Additionally, contributions from gas-phase hydrocarbon photochemistry—possibly seeded by grain-surface products—cannot be ruled out. Further chemical modeling is needed to confirm the efficiency of these pathways for the studied species, which could have important implications for other cold, UV-irradiated environments such as protoplanetary disks.
@@ 1 Introduction
Molecular line observations provide a powerful tool for probing the structure and physical conditions of the molecular gas across a wide range of environments, from high-redshift galaxies (Carilli and Walter 2013) to protoplanetary disks (Öberg et al. 2023). To date, about 330 molecules have been detected in the interstellar medium (ISM) and in circumstellar shells [^1] https://cdms.astro.uni-koeln.de/classic/molecules. While most detections involve simple diatomic or triatomic species, a significant subset consists of organic molecules with six or more atoms, loosely defined as complex organic molecules (COMs; Herbst and van Dishoeck 2009).
@@
The origin of COMs remains an active area of investigation in astrochemistry. Traditionally, COMs have been abundantly detected in hot cores around high-mass protostars (e.g., Bisschop et al. 2007; Feng et al. 2015; Bonfand et al. 2017; van der Walt et al. 2021) and in their low-mass counterparts, hot corinos (e.g., Cazaux et al. 2003; Öberg et al. 2014; Belloche et al. 2020; Yang et al. 2021). Their formation was typically attributed to small organics produced via grain-surface reactions from inherited ices, which are then thermally desorbed and processed further through gas-phase chemistry (Herbst and van Dishoeck 2009). However, the detection of abundant COMs in pre-stellar cores has challenged this view (e.g., Bacmann et al. 2012; Cernicharo et al. 2012; Jiménez-Serra et al. 2016; Scibelli et al. 2024), highlighting the possible importance of cold gas-phase reactions (Shannon et al. 2013), non-diffusive chemistry on grains (Jin and Garrod 2020; Garrod et al. 2022) and non-thermal desorption processes such as chemical desorption (Garrod et al. 2007).
@@
Thus, there is general consensus that the observed abundances of COMs likely arise from an intricate interaction between gas-phase and grain-surface chemistry (Jørgensen et al. 2020; Ceccarelli et al. 2023), though many aspects remain poorly understood. In particular, the role of ultraviolet (UV) radiation is crucial, as it can further process interstellar ices and promote their release into the gas phase through non-thermal photodesorption (Öberg 2016). Moreover, laboratory experiments demonstrate that UV irradiation of ices can even lead to the formation of prebiotic species (de Marcellus et al. 2015; Oba et al. 2019). These processes are especially relevant in environments such as protoplanetary disks, where compelling evidence for in situ ice processing has recently been proposed (e.g., Walsh et al. 2016; Calahan et al. 2023; Ligterink et al. 2024; Yamato et al. 2024; Evans et al. 2025). However, the generally low abundance of COMs and the complex physicochemical structure of these environments (see Öberg et al. 2023, for a recent review) make their study especially challenging.
@@
A promising way to investigate the impact of UV radiation on COM chemistry is to study the edges of molecular clouds illuminated by massive stars, the so-called photodissociation regions (PDRs; Hollenbach and Tielens 1999). In these environments, far-UV (FUV; $6<h\nu<13.6$ eV) photons dominate the chemistry and were initially thought to efficiently photodissociate large molecules. However, more recent studies have challenged this view (see Wolfire et al. 2022, for a recent review). Particularly, COMs have been observed to be abundant in PDRs (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), even when the radiation field exceeds the interstellar average by several orders of magnitude ($G_{0}\sim 10^{2}-10^{4}$, where $G_{0}=1.7$ is the average interstellar radiation field; Draine 1978). Pure gas-phase models cannot account for the observed high COM abundances in these regions (e.g., Gratier et al. 2013; Guzmán et al. 2013; Cuadrado et al. 2017), pointing to a key role for FUV processing of ices. However, the details of this chemistry remain poorly constrained observationally: only a handful of targeted studies exist (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), all relying on single-dish telescopes and thus lacking spatial resolution to determine how COMs are distributed relative to the UV source.
@@
Table 1: Observational parameters for the ALMA-7m+30m maps shown in Figs. 2 and 8.222The projection center of the maps is $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$
@@
Illuminated by the O9.5V binary system $\sigma$ Orionis (Warren and Hesser 1977) $\sim$$3.5$ pc away from the molecular cloud (Abergel et al. 2003), the PDR located at the edge of the Horsehead nebula is particularly well-suited for a spatially resolved study. Thanks to its proximity ($\sim$$400$ pc, Anthony-Twarog 1982), nearly edge-on geometry, and moderate radiation field ($G_{0}\sim 100$, Habart et al. 2005), the Horsehead PDR is an archetype for the average UV radiation fields in molecular clouds of the Milky Way and normal star-forming galaxies. As a consequence, its gas and dust content have been thoroughly characterized at different wavelengths (Schirmer et al. 2020; Hernández-Vera et al. 2023; Abergel et al. 2024, and references therein). Specifically, the Horsehead WHISPER (Wideband High-resolution IRAM-30m Surveys at two Positions with EMIR Receivers, PI: J. Pety) single-dish line survey provided a deep chemical inventory at two key locations [^3] The PDR is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}53\hbox{$.\!\!^{\mathrm{s}}$}936$, $\delta_{2000}=-02^{\circ}28^{\prime}00^{\prime\prime}$ and the dense core is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}55\hbox{$.\!\!^{\mathrm{s}}$}61$, $\delta_{2000}=-02^{\circ}27^{\prime}38^{\prime\prime}$.: the HCO emission peak (PDR position, Gerin et al. 2009) and the DCO+ emission peak (dense core position, Pety et al. 2007). Strikingly, this survey yielded the first PDR detections of complex species such as HCOOH, CH2CO, CH3CHO, and CH3CCH (Guzmán et al. 2014). Since then, several smaller molecules have been mapped at higher angular resolution with the Plateau de Bure Interferometer (PdBI; e.g., Guzmán et al. 2013, 2015) and the Atacama Large Millimeter/submillimeter Array (ALMA; e.g., Hernández-Vera et al. 2023), though similar mapping of COMs remains scarce.
In the present study, we investigate the spatial distribution of O-bearing and N-bearing organics at the UV-illuminated edge of the Horsehead nebula, combining ALMA and IRAM 30m data to achieve $\sim$$15^{\prime\prime}$ resolution (see Fig. 1 for an overview). We examine how gas physical conditions vary across the cloud and explore the chemical formation pathways of COMs under moderate FUV irradiation. Section 2 describes the observations and data combination. Section 3 presents the spatial distributions, physical conditions, and column densities of the analyzed species. In Sect. 4, we elaborate on the possible formation pathways of the studied molecules and discuss implications for FUV-driven organic chemistry in analogous environments. The main results and conclusions are summarized in Sect. 5.
2 Observations and data reduction
2.1 ALMA-7m observations
We used the ALMA Atacama Compact Array (hereinafter ALMA-7m) to perform deep observations of the Horsehead nebula at $\sim 15^{\prime\prime}$ between August 1 2017 and October 1 2017, during Cycle 4 (2016.2.00027.S, PI: V.V. Guzmán). There are 22 successful execution blocks (EBs), each lasting for about $1.5$ h, with an on-source time of $0.8$ h. Single pointing observations centered at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$ were carried out in Band 3 with projected baseline lengths between 5 and 49 m for an average of 10 antennas per EB. The primary beamwidth at $100$ GHz is $107.9^{\prime\prime}$ (see Fig. 1, for a reference in physical units, assuming a distance of $400$ pc). The quasars J0510++1800 or J0522-3627 were observed as the bandpass calibrator; J0423-0120, J0522-3627, or Uranus as the absolute flux calibrator; and J0541-0211 as the phase calibrator.
Due to the small aggregate bandwidth, the bandwidth-switching technique was adopted in order to improve the signal-to-noise ratios of the flux and gain calibration. Another spectral setup was defined for the calibrator scans. The calibration setup includes four wide-band spectral windows in the time division mode (TDM), centering at roughly $98.32$, $100.28$, $110.32$, and $112.32$ GHz with the channel spacing of $15.625$ MHz and the bandwidth of $2.0$ GHz. In each EB, the bandpass calibrator [^4] This is now referred to as the differential gain calibrator and is observed with the DIFFGAIN scan intent in recent cycles. was observed in both FDM (narrow-band) and TDM (wide-band) setups to determine the phase differences between the FDM and TDM windows (see ALMA Technical Handbook for details, e.g. Cortes et al. 2025), while the flux and phase calibrators were observed with the TDM setup only.
The data were initially calibrated manually at the Joint ALMA Observatory (JAO) using the Common Astronomy Software Application (CASA) [^5] https://casa.nrao.edu/ version 5.1.1 (McMullin et al. 2007; CASA Team et al. 2022) with the calibration scripts created by the ALMA Script Generator (Petry et al. 2014). We have modified the scripts and re-calibrated the data with CASA version 5.4.0 in order to optimize the spectral window mapping schemes for the system temperature ($T_{\rm sys}$) and gain (amplitude and phase) calibrations. The $T_{\rm sys}$ of the calibrators and science target data were derived from the measurement scans of the same source and spectral window. The gain solutions for the narrow-band windows of the target were derived from the wide-band windows of the phase calibrator in the same baseband (see Table 2). In addition, antenna CM06 was manually flagged due to unreliable $T_{\rm sys}$ values. After flagging, the on-source visibilities amount to about 14 hours of integration times with 10 antennas.
Figure 2: Zeroth-moment maps gallery for different molecular species moderately resolved at the Horsehead edge. Maps have been rotated $14^{\circ}$ counterclockwise to bring the illuminating star direction in the horizontal direction. Contours are $3,5,10,15,20,25,30]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The red vertical line represents the horizontal zero, delineating the PDR edge (Pety et al. 2005), whereas the cyan and green crosses show the dense core (Pety et al. 2007) and PDR (Gerin et al. 2009) positions, respectively. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Transitions corresponding to the same molecular species are shown using the same color-scale range.
@@ 2.2 IRAM 30m observations
As part of project 108-17, we used the EMIR receivers of the IRAM 30m telescope to map during 47 hours a region of $180^{\prime\prime}\times 180^{\prime\prime}$ centered on ALMA-7m pointing and rotated by $14^{\circ}$ counter-clockwise compared to the Equatorial $2000$ coordinate frame. We used the narrow mode of the Fourier Transform Spectrometer that delivers four times $1.8$ GHz of continuous bandwidths (2 of them in each sideband) at a channel spacing of $50$ kHz. We interleaved two different receiver tunings to cover continuously from the lower sideband from $93.65$ to $100.55$ GHz, and the upper sideband from $109.65$ to $116.65$ GHz. We used the position-switch, on-the-fly observing mode, covering each field with back-and-forth scans along either the right ascension or declination axes. We slewed at a speed of $8.5^{\prime\prime}$ s-1 and we dumped data to disk every $0.5$ s, yielding at least 5 integrations per beam in the scanning direction. The scan legs were separated by $8.5^{\prime\prime}$, yielding Nyquist sampling transverse to the scan direction at $2.6$ mm. We calibrated the system every $15$ minutes.
@@
The calibration of the data was done inside the GILDAS/MRTCAL software and the remaining of the processing happened in the GILDAS/CLASS software. Around the frequency of each line of interest, we extracted a frequency range of $100$ MHz. We baselined the spectra by fitting a polynomial of order 1 after excluding the [$8.5$, $12.5$ km s-1] velocity range. We converted the spectra intensity in main beam temperature using the standard Ruze formula for the IRAM 30m telescope. Finally we resampled the velocity axis to match the ALMA-7m one, and we gridded the data in a cube of $4^{\prime\prime}\times 4^{\prime\prime}$ pixels.

### v3

@@
Complex organic molecules (COMs) are considered essential precursors to prebiotic species in the interstellar and circumstellar medium. Despite their astrobiological relevance, many aspects of the formation of COMs remain unclear, particularly the role of ultraviolet (UV) radiation. While COMs were once expected to be efficiently destroyed under UV-irradiated conditions, detections in photodissociation regions (PDRs) have challenged this view. However, the mechanisms by which UV radiation contributes to their formation are still uncertain. Here we present moderately resolved maps of simple and complex organic molecules at the UV-illuminated edge of the Horsehead nebula, obtained by combining Atacama Large Millimeter/submillimeter Array (ALMA) and IRAM 30m single-dish observations at $\sim$$15^{\prime\prime}$ resolution. For the first time in this PDR environment, we analyzed the spatial distribution of species such as C17O, CH2CO, CH3CHO, HNCO, CH3CN, and HC3N. By incorporating previous C17O and C18O single-dish data as well as Plateau de Bure Interferometer (PdBI) maps of H2CO and CH3OH, we derived profiles of gas density, temperature, thermal pressure, and column densities of the organic species as a function of distance from the UV source. Our results show that most organic species—particularly H2CO, CH2CO, CH3CHO, HNCO, and CH3CN—exhibit enhanced column densities at the UV-illuminated edge compared to cloud interiors, possibly indicating efficient dust-grain surface chemistry driven by the diffusion of atomic C and radicals produced via photodissociation of CO and CH3OH, as supported by recent laboratory experiments. The exceptions, HC3N and CH3OH, can be attributed to inefficient formation on dust grains and ineffective nonthermal desorption into the gas phase, respectively. Additionally, contributions from gas-phase hydrocarbon photochemistry, possibly seeded by grain-surface products, cannot be ruled out. Further chemical modeling is needed to confirm the efficiency of these pathways for the studied species, which could have important implications for other cold UV-irradiated environments such as protoplanetary disks.
@@ 1 Introduction
Molecular line observations provide a powerful tool for probing the structure and physical conditions of the molecular gas across a wide range of environments, from high-redshift galaxies (Carilli and Walter 2013) to protoplanetary disks (Öberg et al. 2023). To date, about 330 molecules have been detected in the interstellar medium (ISM) and in circumstellar shells. [^1] https://cdms.astro.uni-koeln.de/classic/molecules While most detections involve simple diatomic or triatomic species, a significant subset consists of organic molecules with six or more atoms, loosely defined as complex organic molecules (COMs; Herbst and van Dishoeck 2009).
@@
The origin of COMs remains an active area of investigation in astrochemistry. Traditionally, COMs have been abundantly detected in hot cores around high-mass protostars (e.g., Bisschop et al. 2007; Feng et al. 2015; Bonfand et al. 2017; van der Walt et al. 2021) and in their low-mass counterparts, hot corinos (e.g., Cazaux et al. 2003; Öberg et al. 2014; Belloche et al. 2020; Yang et al. 2021). Their formation was typically attributed to small organics produced via grain-surface reactions from inherited ices, which are then thermally desorbed and processed further through gas-phase chemistry (Herbst and van Dishoeck 2009). However, the detection of abundant COMs in pre-stellar cores has challenged this view (e.g., Bacmann et al. 2012; Cernicharo et al. 2012; Jiménez-Serra et al. 2016; Scibelli et al. 2024), highlighting the possible importance of cold gas-phase reactions (Shannon et al. 2013), nondiffusive chemistry on grains (Jin and Garrod 2020; Garrod et al. 2022), and nonthermal desorption processes such as chemical desorption (Garrod et al. 2007).
@@
Thus, there is general consensus that the observed abundances of COMs likely arise from an intricate interaction between gas-phase and grain-surface chemistry (Jørgensen et al. 2020; Ceccarelli et al. 2023), though many aspects remain poorly understood. In particular, the role of ultraviolet (UV) radiation is crucial as it can further process interstellar ices and promote their release into the gas phase through nonthermal photodesorption (Öberg 2016). Moreover, laboratory experiments demonstrate that UV irradiation of ices can even lead to the formation of prebiotic species (de Marcellus et al. 2015; Oba et al. 2019). These processes are especially relevant in environments such as protoplanetary disks, where compelling evidence for in situ ice processing has recently been proposed (e.g., Walsh et al. 2016; Calahan et al. 2023; Ligterink et al. 2024; Yamato et al. 2024; Evans et al. 2025). However, the generally low abundance of COMs and the complex physicochemical structure of these environments (see Öberg et al. 2023, for a recent review) make their study especially challenging.
@@
A promising way to investigate the impact of UV radiation on COM chemistry is to study the edges of molecular clouds illuminated by massive stars, known as the photodissociation regions (PDRs; Hollenbach and Tielens 1999). In these environments, far-UV (FUV; $6<h\nu<13.6$ eV) photons dominate the chemistry and were initially thought to efficiently photodissociate large molecules. However, more recent studies have challenged this view (see Wolfire et al. 2022, for a recent review). Particularly, COMs have been observed to be abundant in PDRs (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), even when the radiation field exceeds the interstellar average by several orders of magnitude ($G_{0}\sim 10^{2}-10^{4}$, where $G_{0}=1.7$ is the average interstellar radiation field; Draine 1978). Pure gas-phase models cannot account for the observed high COM abundances in these regions (e.g., Gratier et al. 2013; Guzmán et al. 2013; Cuadrado et al. 2017), pointing to a key role for FUV processing of ices. However, the details of this chemistry remain poorly constrained observationally: only a handful of targeted studies exist (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), all relying on single-dish telescopes, and thus lacking spatial resolution to determine how COMs are distributed relative to the UV source.
Illuminated by the O9.5V binary system $\sigma$ Orionis (Warren and Hesser 1977) $\sim$$3.5$ pc away from the molecular cloud (Abergel et al. 2003), the PDR located at the edge of the Horsehead nebula is particularly well-suited for a spatially resolved study. Thanks to its proximity ($\sim$$400$ pc, Anthony-Twarog 1982), nearly edge-on geometry, and moderate radiation field ($G_{0}\sim 100$, Habart et al. 2005), the Horsehead PDR is an archetype for the average UV radiation fields in molecular clouds of the Milky Way and normal star-forming galaxies. As a consequence, its gas and dust content have been thoroughly characterized at different wavelengths (Schirmer et al. 2020; Hernández-Vera et al. 2023; Abergel et al. 2024, and references therein). Specifically, the Horsehead Wideband High-resolution IRAM 30m Surveys at two Positions with EMIR Receivers (WHISPER, PI: J. Pety) single-dish line survey provided a deep chemical inventory at two key locations: [^2] The PDR is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}53\hbox{$.\!\!^{\mathrm{s}}$}936$, $\delta_{2000}=-02^{\circ}28^{\prime}00^{\prime\prime}$ and the dense core is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}55\hbox{$.\!\!^{\mathrm{s}}$}61$, $\delta_{2000}=-02^{\circ}27^{\prime}38^{\prime\prime}$. the HCO emission peak (PDR position, Gerin et al. 2009) and the DCO+ emission peak (dense core position, Pety et al. 2007). Strikingly, this survey yielded the first PDR detections of complex species such as HCOOH, CH2CO, CH3CHO, and CH3CCH (Guzmán et al. 2014). Since then, several smaller molecules have been mapped at higher angular resolution with the Plateau de Bure Interferometer (PdBI; e.g., Guzmán et al. 2013, 2015) and the Atacama Large Millimeter/submillimeter Array (ALMA; e.g., Hernández-Vera et al. 2023), though similar mapping of COMs remains scarce.
For the present study, we investigated the spatial distribution of O-bearing and N-bearing organics at the UV-illuminated edge of the Horsehead nebula, combining ALMA and IRAM 30m data to achieve $\sim$$15^{\prime\prime}$ resolution (see Fig. 1 for an overview). We examined how the physical conditions of gas vary across the cloud and explored the chemical formation pathways of COMs under moderate FUV irradiation. Section 2 describes the observations and data combination. Section 3 presents the spatial distributions, physical conditions, and column densities of the analyzed species. In Sect. 4 we elaborate on the possible formation pathways of the studied molecules and discuss implications for FUV-driven organic chemistry in analogous environments. The main results and our conclusions are summarized in Sect. 5.
2 Observations and data reduction
2.1 ALMA 7m observations
We used the ALMA Atacama Compact Array (hereafter ALMA 7m) to perform deep observations of the Horsehead nebula at $\sim 15^{\prime\prime}$ between August 1, 2017, and October 1, 2017, during Cycle 4 (2016.2.00027.S, PI: V.V. Guzmán). There were 22 successful execution blocks (EBs), each lasting for about $1.5$ h, with an on-source time of $0.8$ h. Single pointing observations centered at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$ were carried out in Band 3 with projected baseline lengths between 5 and 49 m for an average of ten antennas per EB. The primary beamwidth at $100$ GHz is $107.9^{\prime\prime}$ (see Fig. 1, for a reference in physical units, assuming a distance of $400$ pc). Quasar J0510++1800 or J0522-3627 was observed as the bandpass calibrator; J0423-0120, J0522-3627, or Uranus as the absolute flux calibrator; and J0541-0211 as the phase calibrator.
Due to the small aggregate bandwidth, the bandwidth-switching technique was adopted in order to improve the signal-to-noise ratios of the flux and gain calibration. Another spectral setup was defined for the calibrator scans. The calibration setup includes four wide-band spectral windows in the time division mode (TDM), centering at roughly $98.32$, $100.28$, $110.32$, and $112.32$ GHz with the channel spacing of $15.625$ MHz and the bandwidth of $2.0$ GHz. In each EB, the bandpass calibrator [^3] This is now referred to as the differential gain calibrator and is observed with the DIFFGAIN scan intent in recent cycles. was observed in both FDM (narrow-band) and TDM (wide-band) setups to determine the phase differences between the FDM and TDM windows (see, e.g., ALMA Technical Handbook for details Cortes et al. 2025), while the flux and phase calibrators were observed with the TDM setup only.
The data were initially calibrated manually at the Joint ALMA Observatory (JAO) using the Common Astronomy Software Application (CASA) [^4] https://casa.nrao.edu/ version 5.1.1 (McMullin et al. 2007; CASA Team et al. 2022) with the calibration scripts created by the ALMA Script Generator (Petry et al. 2014). We have modified the scripts and recalibrated the data with CASA version 5.4.0 in order to optimize the spectral window mapping schemes for the system temperature ($T_{\rm sys}$) and gain (amplitude and phase) calibrations. The $T_{\rm sys}$ of the calibrators and science target data were derived from the measurement scans of the same source and spectral window. The gain solutions for the narrow-band windows of the target were derived from the wide-band windows of the phase calibrator in the same baseband (see Table 2). In addition, antenna CM06 was manually flagged due to unreliable $T_{\rm sys}$ values. After flagging, the on-source visibilities amount to about 14 hours of integration times with ten antennas.
@@ 2.1 ALMA 7m observations
Table 1: Observational parameters for the ALMA 7m+30m maps shown in Figs. 2 and 8.555The projection center of the maps is $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$
@@ 2.2 IRAM 30m observations
As part of project 108-17, we used the EMIR receivers of the IRAM 30m telescope to map during 47 hours a region of $180^{\prime\prime}\times 180^{\prime\prime}$ centered on ALMA 7m pointing and rotated by $14^{\circ}$ counterclockwise compared to the Equatorial $2000$ coordinate frame. We used the narrow mode of the Fourier Transform Spectrometer that delivers four times $1.8$ GHz of continuous bandwidths (two of them in each sideband) at a channel spacing of $50$ kHz. We interleaved two different receiver tunings to cover continuously from the lower sideband from $93.65$ to $100.55$ GHz, and the upper sideband from $109.65$ to $116.65$ GHz. We used the position-switch, on-the-fly observing mode, covering each field with back-and-forth scans along either the right ascension or declination axes. We slewed at a speed of $8.5^{\prime\prime}$ s-1 and we dumped data to disk every $0.5$ s, yielding at least 5 integrations per beam in the scanning direction. The scan legs were separated by $8.5^{\prime\prime}$, yielding Nyquist sampling transverse to the scan direction at $2.6$ mm. We calibrated the system every $15$ minutes.
@@
The calibration of the data was done inside the GILDAS/MRTCAL software and the remaining of the processing happened in the GILDAS/CLASS software. Around the frequency of each line of interest, we extracted a frequency range of $100$ MHz. We baselined the spectra by fitting a polynomial of order 1 after excluding the [$8.5$, $12.5$ km s-1] velocity range. We converted the spectra intensity in main beam temperature using the standard Ruze formula for the IRAM 30m telescope. Finally we resampled the velocity axis to match the ALMA 7m axis, and we gridded the data in a cube of $4^{\prime\prime}\times 4^{\prime\prime}$ pixels.
2.3 Interferometric and single-dish data combination
We combined the IRAM 30m and ALMA 7m to produce a clean image containing all spatial information from 0 to $50$ m, i.e., equivalent to a $50$ m diameter total-power telescope. The IRAM 30m map is deconvolved from the $30$ m beam in the Fourier plane before multiplication by the ALMA 7m primary beam in the image plane. After a last Fourier transform, pseudo-visibilities were sampled between 0 and $7$ m. These visibilities were then merged with the ALMA 7m interferometric observations (see Rodriguez-Fernandez et al. 2008, for details). Standard imaging, CLEAN deconvolution, and primary beam correction steps were then applied. The resulting data cube was then scaled from Jy beam-1 to $T_{\mathrm{mb}}$ temperature scale using the synthesized beam size. All these processes happened in the GILDAS/MAPPING software.
The observational parameters of the final combined (hereafter ALMA 7m+30m) images are summarized in Table 1. The channel rms is calculated using the central 50% of the pixels in the first and last five (line-free) channels of the data cubes. To compare the spatial distribution of the different molecular tracers, we collapse the line emission into velocity-integrated intensity, or zeroth-moment, maps using the CASA task immoments. The zeroth-moment rms is estimated as $\sigma_{\mathrm{rms}}\approx\sigma_{\mathrm{chan}}\sqrt{N_{\mathrm{chan}}}\Delta v$, where $\sigma_{\mathrm{chan}}$ is the channel rms, $N_{\mathrm{chan}}$ is the number of channels used to compute the zeroth-moment, and $\Delta v$ is the channel spacing in velocity.
2.4 Legacy data
The analysis of the ALMA 7m+30m data was complemented with PdBI observations of formaldehyde (H2CO) $J_{K_{a},K_{c}}=2_{02}-1_{01}$ and methanol (CH3OH) $J_{K}=3_{-1}-2_{-1}$ (E symmetry) rotational transitions imaged at $\theta_{\mathrm{beam}}\sim 6^{\prime\prime}$. All the observational details of these legacy data can be found in Guzmán et al. (2013). In order to make the two data sets comparable, the PdBI data was rotated by 14, and the angular resolution was downgraded to $\theta_{\mathrm{beam}}\sim 15^{\prime\prime}$ to match the ALMA 7m+30m data using the CASA tools rotate and convolve2d, respectively. Afterward, the spatial axes were regridded using the CASA task imregrid to match the number of pixels per beam. Throughout this procedure, the CH3CN image was used as a template. Thus, for all subsequent analysis, the angular resolution of the H2CO and CH3OH images is exactly the same as for CH3CN (see Table 1).
Figure 2: Zeroth-moment maps gallery for different molecular species moderately resolved at the Horsehead edge. The maps have been rotated $14^{\circ}$ counterclockwise to bring the illuminating star direction in the horizontal direction. The contours are $3,5,10,15,20,25,30]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The red vertical line represents the horizontal zero, delineating the PDR edge (Pety et al. 2005), whereas the cyan and green crosses show the dense core (Pety et al. 2007) and PDR (Gerin et al. 2009) positions, respectively. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Transitions corresponding to the same molecular species are shown using the same color-scale range.